A man runs diagonally across from one corner of a rectangular field 80m by 60m to the opposite corner in a straight line speed of 9 meters per second.find the time taken for him to complete the run.
Answer by rfer(16322)

60^2+80^2=c^2
c=sqrt 10000
c=100
100/9=11.1 sec